OpenStudy (anonymous):

Fooool's problems of the day,two *very very* easy problem: 1) How many four digit numbers of the form AABA are divisible by \(33\)? 2) How many three digit even numbers are there such that if they are subtracted from their reverse the result is a positive multiple of \(66\)? Lets see, who can come up with the solution first!

OpenStudy (anonymous):

1. 0?

OpenStudy (king):

no i found one for the 1st one its not 0....wait...

OpenStudy (anonymous):

I am sorry Ishaan but that's not the right answer.

OpenStudy (anonymous):

\[\frac{{AA*100} + BA}{3*11} = \frac{AA*100}{11*3} + \frac{BA}{11*3}\] I don't think there is a number in the form of BA divisible by 11. BTW What do you mean by AABA, numbers like 1121, 1131, 2242?

OpenStudy (anonymous):

\( 3333 \) is a valid isn't? ;)

OpenStudy (anonymous):

But you said AABA? Hmm if B=A is valid then, {3333, 6666, 9999} are clear solutions.

OpenStudy (anonymous):

I never said \(B \neq A \) ;)

OpenStudy (mr.math):

1) All numbers in the form \(AABA\) such that: i) \(3A+B=k\), where k is a number divisible by 3, and ii) \(B-A=c\), where c is divisible by 11. Now, for k, we have \(0\le k \le 36\). There are \(\frac{36}{3}+1=13\) such k in this interval. [That's {0,3,6,..36}] For c, we have \(0\le c \le 9\), there's only one c that's divisible by 11 which c=0. Hence \[B-A=0 \implies B=A\] and \[3A+B=0, 3, 6,.. 36\] From here we can conclude that the numbers are {0000, 3333, 6666, 9999}.

OpenStudy (king):

answer is 3?

OpenStudy (mr.math):

4.

OpenStudy (king):

0000 is 0 not a 4 digit no. ...........

OpenStudy (anonymous):

Hmm then Zero? //Now it's confusing me, I am moving on to Second part. Let the three Digit even number be ABC ABC - CBA Now for even ABC, C has to even A*100 + B*10 + C - C*100 -B*10 -A = 66k A(100 - 1) - C (100- 1) = 66k 3*33(A-C) = 2*33*k \[\frac{3}{2} * (A-C) = k\] We know C is even and to satisfy k as an Integer A has to be even too (two in the denominator) So, all the numbers in the form of ABC which have A and C as even numbers.

OpenStudy (mr.math):

Hmm, I need to check the definition of a digit. But the idea is clear, I think.

OpenStudy (king):

@ishaan dont we have to find the amnt?

OpenStudy (anonymous):

0000 is not a four digit number.

OpenStudy (king):

yes!!!

OpenStudy (mr.math):

Then 3 such numbers exist.

OpenStudy (anonymous):

Try the second one now :)

OpenStudy (king):

i fnd 1 number fr the 2nd one..... 400

OpenStudy (mr.math):

I think Ishaan already did it.

OpenStudy (king):

but then we hav o find the amnt

OpenStudy (king):

is 0 div by 66?

OpenStudy (mr.math):

I would rather let him find it, or I would be stealing his work :P

OpenStudy (anonymous):

Ishaan's answer is probably 4*10*5 but that doesn't seem to be the correct answer.

OpenStudy (king):

is 0 div. by 66.....

OpenStudy (anonymous):

Oh, I wasn't trying to find the number, Limit of laziness. Let me try now!

OpenStudy (nottim):

How do you learn how to solve these problems??

OpenStudy (king):

is the number 202,or 404 or 101 counted ?

OpenStudy (mr.math):

I would add to Ishaan solution that \[100A+10B+C\ge 100C+10B+A \implies A\ge C.\]

OpenStudy (mr.math):

This is probably the only missing piece in his solution.

OpenStudy (king):

guys are palindromic nos. counted?

OpenStudy (king):

palindromic even nos.?

OpenStudy (anonymous):

Hmm according to me I need even numbers for A and C of the form ABC. \[A \in {2,4,6,8}\] Oh yeah, Mr.Math is right.

OpenStudy (mr.math):

Oh or \(A>C\), since you said positive (i.e. not including difference of 0).

OpenStudy (king):

so, ishaan is 202,404,606,808 rite?

OpenStudy (anonymous):

Then the number is certainly less than 200. No@ King, the question says positive multiples. The number must be for 200s 0. For 400s 9 (412, 422, 432, ...). For 600s 18 (I think not sure) and finally for 800s 27. Is it 9 + 18 + 27 = 9(1 + 2+3) = 54?

OpenStudy (anonymous):

The answer is 60.

OpenStudy (anonymous):

Oh no :-/ I must have done something silly.

OpenStudy (king):

there are 21 in 600's and 30 in 800's so answer is 60

OpenStudy (anonymous):

Oh yeah, How can I forget 802, 804, 806, 604, 602 and 402!?!

OpenStudy (anonymous):

Well done ishaan :)

OpenStudy (anonymous):

Thanks! Took me 40 minutes to solve a very very easy problems of yours, I wonder how much hard one would take.

OpenStudy (anonymous):

hehe :)